Operations Report – February 20th

Crew 189 Operations Report 20Feb2018
SOL: 2

Name of person filing report: PAYEN Gabriel

Non-nominal systems: Backpack 3


Solar panels < 70% without generator running

Notes on non-nominal systems:

Backpack 3 :

All engineering checks were nominals, but during the EVA both fans turned out of service. Also, the necklace which holds the helmet was not stretched enough so that the helmet fell on the neck of Gabriel during half of the EVA.

Backpack 3 joined backpack 4 to the list of backpacks the crew need to fix. We’ll certainly begin to fix them tomorrow.

WiFi :

Only 15 Mo were left this morning (immediately after the very connection of the day were established). Yet, the leak is unknow. I reported the problem this morning and the crew have communicated with CapCom to find a solution.

Solar panels :

I had to go out of the SIM to turn on the generator at 6:20 PM. Solar SOC was at 53%, I hope batteries are not injured.

Generator (hours run): Turned of at 12:00 and turned on at 6:20PM.

Solar— SOC % (Before generator is run at night) /!\ 53%

Diesel – 70%

Propane – 40 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 7 gallons.

Water (trailer) – 0 gallons.

Water (static) – 65% gallons

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 0.26m

Toilet tank emptied: NO

ATV’s Used: NO

Deimos rover used: YES

Hours: 109.1

Beginning charge: 100%

Ending charge: 100% (the EVA team doesn’t travel a lot)

Currently charging: YES

Spirit rover used: NO

Hours: 20.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: NO

Opportunity rover used: YES

Hours: 16.6

Beginning charge: 100%

Ending charge: 45%

Currently charging: YES

Curiosity rover used: NO

Hours: 100%

Beginning charge: –

Ending charge: –

Currently charging: NO

General notes and comments: –

Summary of internet:

-Still seeking the leak.

Summary of suits and radios:

-Backpack 3 is put aside. Starting fixing backpack 3 and 4 tomorrow.

Summary of Hab operations:

-Two thermometers and the shape for the GreenHab received in the airlock. Jeremy (GreenHab officer) will set it up tomorrow.

-The pipe in the low desk seems not to flow anymore. All water devices in the Hab and GreenHab were tested one by one and none trigger the leak.

-Considering a Dr Rupert’s request, a mattress will be put in the airlock between 8PM and 9 PM.

Summary of GreenHab operations: –

Summary of ScienceDome operations: –

Summary of RAMM operations: –

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support: –

EVA Report – February 20th

EVA Report:

EVA #2 Crew members: Victoria Da-Poian (EVA leader), Jérémy Auclair, Gabriel Payen, Benoit Floquet

Location: – Around the Hab 12S-518100E-4250700N

Time: departure at 09:30 a.m. Duration: 2 hour 40 minutes

Narration: I was really happy to explore again the red cliffs with my new crew. Since this morning, everyone has experienced the EVAs. We managed to deploy one experiment (the MegaARES) but a small problem interrupted the LOAC deployment. We will fix it with Jérémy in order to try again tomorrow. Even if Gabriel had a very foggy helmet (fans problems), and everyone seems really happy and I have to say the photos Benoit took are really nice. During our EVA, we had a very nice teamwork. I am very proud of our common work and I am looking forward to going back to EVA with them.

Sol Summary – February 20th

Crew 189 Sol Summary Report 20 Feb 2018

Sol 2

Summary Title: First experiments deployment

Author: Victoria Da-Poian

Mission Status: All nominal

Sol Activity Summary: Our day began with our lovely physical training.
It is already adopted by each crewmember. Since this morning, everyone
has experienced the EVAs. We managed to deploy one experiment (the
MegaARES) but a small problem enabled the LOAC deployment. We will fix
it with Jérémy in order to try again tomorrow. Even if Gabriel had a
very foggy helmet (fans problems), and everyone seems really happy and
I have to say the photos Benoit took are really nice. During our EVA,
the rest of the crew fixed the water boiler, cleaned, cooked and
worked on experiments. After the lunch and a little break, Gabriel
showed us his humans factors experiment, and everyone worked on their
own experiments and reports.

Look Ahead Plan: After our first deployment, the second air sensor,
(LOAC) will normally be deployed tomorrow.

Anomalies in work: Water leak in the Hab Lower deck, broken fans in spacesuits

Weather: Nice

Crew Physical Status: Fine, already a bit tired from the first EVAs

Reports to be filed
Greenhab Officer

Support Requested: No more boiler is needed. Would it be possible to
have the RoverCal that is in the RAM delivered in the airlock please?

Greenhab Report – February 20th

Jérémy Auclair

20 February 2018

Environmental control:

Ambient with door open

Fan during 30 minutes

Shade cloth on

Average temperatures: I received the thermometers only tonight, I will use it tomorrow. Nevertheless I estimated the temperature around 1pm , about 30 °C.

Hours of supplemental light: from 6pm to 11pm

Changes to crops: Three cherry tomatoes harvested

Daily water usage for crops: 13 gallons

Time(s) of watering for crops: 12:30pm and 6:45pm

Aquaponics: None

Narrative: I think I will have some work of reorganization of the plants. After all my experiments set I will do that.

Support/supplies needed: Wooden sticks to support different plants

Commander Report – February 20th

Crew 189 Commander Report 20Feb2018

Sol 2

Today was probably our first “normal” day: everything is now almost
set up, excluding our experiments, that are for most of them already
facing some predictable minor issues. It is also the day I started to
monitor our water consumption. Like during our previous rotation, we
keep track of filtered water, greenhab water, cooking water, flushes
and showers, to understand our consumption better.

We woke up at 7:00 am, exercised together, and took extra time to
prepare ourselves better for our second EVA. This one had a bigger
scientific role than yesterday’s one: we wanted to deploy both the
MegaAres and the LOAC near Pooh’s Corner, in the spot we had chosen
yesterday. My fellow commander Victoria who was EVA leader today, went
with Jérémy, Gabriel, Benoit and one of our school’s mascot to the
work place. They first tried to deploy the LOAC experiment, but a wire
connection blocked the deployment. After this first failure, they
worked altogether on the MegaARES. Two hours, 10 nails, and many
screws later, it was operational and they all went back happy of this
EVA, looking forward to deploying the other experiments. Stay tuned
tomorrow to know more about the LOAC deployment.

At the same time in the hab, we cooked, kept assembling Laurent’s
solar panels cleaners, and dissembled the water boiler to try to fix
it. It was only when Benoit came back with the voltmeter that he
managed to fix it with a little soldering. This may seem secondary,
but the whole food needs to be rehydrated using boiling water so that
a working boiler allows us to accelerate every meal preparation. I
then tried to bake bread and ended up messing with cups and US
tablespoon so that we had to rename it “the Rock”…

This afternoon started by a victory from the commander’s duo in our
French card game “coinche” at the end of lunch break. Then, we
experienced several informatic problems while trying to launch our
human factor experiments, but nothing serious. We will just have to
wait for one more day to launch these. The end of the afternoon was
quiet as everybody was busy working on experiments and reports.

Ad Astra!

Journalist Report – February 20th

Journalist report, 02/20/18: First scientific and gastronomic hitches

Our first own made bread prepared by Victoria was welcomed after such a long day we had. Though it was somehow a bit compact, it was yet tasty, and with some strawberry jam it made a nice dessert! After having filled our stomachs, it was yet time to go to sleep, to prepare for the second day of mission.

And sleeping early was crucial, as today’s sport beginning time was 7am … Some of our fellow crew members wore their nicest outfits : Benoît put on his shirt of Paris Saint-Germain, while Gabriel chosed the winning combo : fitted black shirt and (very) fitted flashy yellow legging. However, the sport session was shortened for the ones who would go on EVA: Gabriel, Jérémy, Benoît and Commander Victoria were the ones to go out this morning. The EVA followed the purpose of the last one: it was now time to place the measurement instruments, after having chosen the location, according to our observations of the previous day.

However, everything did not work as we expected. The team took a lot more time to be ready than they thought, as they had a lot of material to bring on the rover. The team first went placing the instrument of Jérémy’s experiment, the LOAC. Though it was perfectly working in the Hab, it did not give any result on the outside. After a small check, the team realized the alimentation was broken, and that they had to bring back the LOAC to the station to fix it. The second instrument, Gabriel’s MegaARES, took almost 2 hours to the team to be placed. Gabriel had indeed some issues with his helmet, which started to be covered by fog: he was blinded, and had thus to give instructions to the others members of the EVA team, which was not an easy thing to do. The team learned that some tasks which were usually easy became nearly impossible wearing a spacesuit: to tighten a lonely screw took the team over half an hour, yet they finally succeeded with the help of a perfect teamwork. At the same time in the Hab, Louis wanted to follow Victoria’s example in baking new bread. However, he realized later that he did not put the ingredients in good quantities (especially because of American cooking units, which Louis was not happy to be forced to use).

After the team got back in the station, Jérémy was happy to find 3 red cherry tomatoes in the GreenHab : our first fresh products from there. Those were not really large (especially the one we named Victoria), yet it did not reduce our satisfaction. The afternoon was slower: Benoît learned to play French game “coinche”, but he and Jérémy lost nonetheless in a large way against Louis and Victoria. As we reached the middle of the afternoon, Louis’ bread was finally ready. Yet, even though it looked at first more like a Martian rock than like a French baguette, it was quite comestible after all. It was then time for Gabriel and Jérémy to launch their human factors experiments. However, they suffered some software issues, and so were not able to make their first tries today: these will so normally begin tomorrow. These missed tries conclude a day full of small issues, yet things will certainly get better as soon as tomorrow!

Alexandre Martin, Crew 189 Journalist

Copyright © The Mars Society. All rights reserved. | Main Site